# FastTrackTr:Towards Fast Multi-Object Tracking with Transformers

### v2

@@ 4.3 Ablation Study
In Table 6, we demonstrate the impact of different trajectory cross-attention mechanisms within the trajectory decoder on tracking performance. Here, $q^{t}$ represents the original query, while $q_{f}^{t-1}$ denotes the output of the decoder from the previous frame, with ’concat’ indicating their concatenation. It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the trajectory encoder, hence its performance is not as good as that of the final model.
@@
Table 7 analyzes the effect of different training frame counts on long-term fusion performance. Utilizing 8 frames yields the highest HOTA, demonstrating superior tracking accuracy. Compared to using fewer frames, 8 frames offer substantial performance improvements while maintaining efficient training times. Beyond 8 frames, metrics show minimal or negligible gains, indicating diminishing returns with additional frames. Thus, selecting 8 training frames provides an optimal balance between enhanced performance and training efficiency for FastTrackTr.
Table 7: Training frames for long-term fusion.
@@ Appendix B How to Improve MOTR inference Speed
This led us to speculate whether the issue might be due to the dynamic nature of the query system, which could result in variable memory usage. As shown in Figure 6, we compare the memory consumption of the decoder when using fixed queries versus dynamic queries. After further investigation, we found that this is indeed a plausible cause. In PyTorch, memory is managed through cudaMalloc and cudaFree, and this dynamic memory management approach could lead to frequent memory allocation calls in models like MOTR, where memory demand fluctuates. This could slow down the model and prevent full GPU utilization.

### v3

@@ 4.3 Ablation Study
In Table 6, we demonstrate the impact of different trajectory cross-attention mechanisms within the trajectory decoder on tracking performance. Here, $q^{t}$ represents the original query, while $q_{f}^{t-1}$ denotes the output of the decoder from the previous frame, with ’concat’ indicating their concatenation. It is important to note that in this table, $q_{f}^{t-1}$ has not passed through the historical encoder, hence its performance is not as good as that of the final model.
@@
Table 7 analyzes the effect of different length of video clips during training. Utilizing 8 frames yields the highest HOTA, demonstrating superior tracking accuracy. Compared to using fewer frames, 8 frames offer substantial performance improvements while maintaining efficient training times. Beyond 8 frames, metrics show minimal or negligible gains, indicating diminishing returns with additional frames. Thus, selecting 8 training frames provides an optimal balance between enhanced performance and training efficiency for FastTrackTr.
Table 7: The length of video clips during training.
@@ Appendix B How to Improve MOTR inference Speed
This led us to speculate whether the issue might be due to the dynamic nature of the query system, which could result in variable memory usage. As shown in Figure 6, we compare the memory consumption of the decoder when using fixed queries versus dynamic queries. After further investigation, we found that this is indeed a plausible cause. In cuda, memory is managed through cudaMalloc and cudaFree. These operations incur significant computational overhead. To mitigate this, PyTorch attempts to reuse memory blocks previously allocated via cudaMalloc. When a suitable cached block exists in PyTorch’s allocator, it directly serves the request without invoking cudaMalloc, thereby minimizing latency during initial allocation phases. However, MOTR’s dynamic data patterns require intermediate tensors of varying sizes throughout runtime. This variability prevents the allocator from leveraging cached blocks of appropriate dimensions, forcing frequent calls to cudaFree to release outdated allocations and subsequent cudaMalloc requests for new memory spaces. Consequently, the allocator repeatedly rebuilds its caching infrastructure through bulk CUDA API operations – an expensive process that introduces substantial latency spikes. This thrashing behavior substantially degrades computational efficiency and hinders full GPU utilization